### v1

# Contextuality from Single-State Representations: An Information-Theoretic Principle for Adaptive Intelligence

Abstract
Adaptive systems often operate across multiple contexts while reusing a fixed internal state space due to constraints on memory, representation, or physical resources. Such single-state reuse is ubiquitous in natural and artificial intelligence, yet its fundamental representational consequences remain poorly understood. We show that contextuality is not a peculiarity of quantum mechanics, but an inevitable consequence of single-state reuse in classical probabilistic representations. Modeling contexts as interventions acting on a shared internal state, we prove that any classical model reproducing contextual outcome statistics must incur an irreducible information-theoretic cost: dependence on context cannot be mediated solely through the internal state. We provide a minimal constructive example that explicitly realizes this cost and clarifies its operational meaning. We further explain how nonclassical probabilistic frameworks avoid this obstruction by relaxing the assumption of a single global joint probability space, without invoking quantum dynamics or Hilbert space structure. Our results identify contextuality as a general representational constraint on adaptive intelligence, independent of physical implementation.
I Introduction
Adaptive systems—biological or artificial—must routinely operate across multiple contexts while relying on limited internal resources. In many realistic settings, such systems are constrained to maintain a single internal state space that is reused across different interactions, measurements, or decision scenarios. Here, “single-state” does not refer to a scalar or binary variable, but to a fixed internal state space—potentially high-dimensional and carrying many bits of information—that cannot be duplicated, indexed, or branched according to context. Such constraints arise naturally in perceptual systems with limited working memory, decision-making agents with bounded internal representations, and physical devices whose internal degrees of freedom must be reused across contexts. Related constraints have long been discussed across cognitive science and artificial intelligence under notions such as bounded rationality, resource-limited agents, and fixed internal representational architectures (e.g., Herbert A. Simon; Models of Bounded Rationality [4]; John R. Anderson [1]). Despite their ubiquity, the representational consequences of maintaining a fixed internal state space across contexts remain poorly understood.
A central difficulty lies in how contextual dependence is represented under such constraints. Standard classical approaches typically address contextuality by introducing external labels, hidden variables, or context-dependent internal states, effectively enlarging the representational space. While successful in many applications, these strategies implicitly relax the single-state constraint by allowing contextual information to be stored or accessed explicitly. In contrast, many adaptive systems—such as perceptual agents responding to sequential stimuli under limited working memory or decision systems operating under strict memory reuse—must accommodate contextual changes through interventions that modify information processing, rather than through explicit encoding of which context is active (e.g., Working Memory [2]). This distinction motivates a fundamental question: what representational and informational costs are unavoidable when a system is required to maintain a fixed internal state space across multiple interventions?
In this work, we address this question at the level of representation rather than dynamics or physical implementation. We consider systems whose internal state space may be high-dimensional, but must remain fixed and internally consistent across contexts. Within this minimal framework, we show that maintaining such representations necessarily entails an irreducible information-theoretic cost when contextual dependence is present. Crucially, this cost does not arise from insufficient state capacity, but from the requirement that a single internal state space be reused coherently across multiple interventions. This perspective reframes contextuality not as a peculiarity of particular physical theories, but as a general and unavoidable constraint on representation.
Our approach makes three complementary contributions. First, we introduce a minimal operational formulation of fixed-state representations that cleanly separates contextual effects from assumptions about dynamics or physical realization. Second, we demonstrate that any classical probabilistic representation satisfying this formulation necessarily incurs an irreducible information-theoretic cost when contextual dependence is present, reflecting a fundamental obstruction to representing multiple interventions within a single internally consistent state space. Third, we present a simple constructive mechanism showing how nonclassical statistical structure can emerge directly from representational constraints, without invoking quantum postulates or Hilbert space assumptions. Together, these results establish contextuality as a representational principle, rather than a feature tied to specific physical theories.
The remainder of this paper is organized as follows. In Sec. II, we introduce a representational framework based on single-state reuse and contextual interventions, and define the class of classical probabilistic representations considered throughout. In Sec. III, we establish the central information-theoretic obstruction, proving that classical single-state representations necessarily incur an irreducible contextual information cost under contextual dependence. Sec. IV provides a minimal constructive illustration showing that this obstruction corresponds to realizable representational behavior rather than a purely formal limitation. In Sec. V, we explain why nonclassical probabilistic frameworks naturally accommodate single-state reuse without additional information cost. Sec. VI discusses the implications of these results for adaptive and intelligent systems operating under resource constraints, and Sec. VII concludes with a discussion of limitations and future directions.
II Representational Framework
II.1 Single-State Representations and Interventions
We begin by formalizing the representational framework considered throughout this work. Our goal is to isolate constraints on representation itself, independently of dynamics, learning rules, or physical implementation.
Definition 1 (Single-State Representation).
A single-state representation consists of a fixed internal state space $\mathcal{S}$ together with a family of response functions that map elements of $\mathcal{S}$ to observable outcomes. The state space $\mathcal{S}$ may be high-dimensional and encode many bits of information, but it is not indexed, duplicated, or partitioned according to context. All contextual variation must therefore be mediated through transformations acting on $\mathcal{S}$, rather than through context-dependent enlargement or branching of the state space.
Importantly, the single-state condition does not impose a bound on state capacity. Instead, it constrains how internal information may be reused across contexts: the same underlying state space must support all contextual interactions.
Definition 2 (Interventions).
A context is modeled as an intervention, defined as a map that acts on the internal state space $\mathcal{S}$ or on the response functions associated with $\mathcal{S}$. Interventions modify how the internal state is processed or probed, but do not introduce additional state variables or external context labels accessible to the representation. In particular, the identity of the intervention is not assumed to be explicitly encoded within $\mathcal{S}$.
This notion distinguishes interventions from classical context labels: while labels index distinct representational resources, interventions act on a common resource that must be reused across contexts.
Definition 3 (Representation Consistency).
A single-state representation is said to be consistent if all interventions act on the same internal state space $\mathcal{S}$ and yield outcome statistics that are jointly representable without contradiction. Consistency requires that the internal representation does not depend on which interventions are applied before or after, except through their defined action on $\mathcal{S}$. In particular, no intervention-specific hidden states or context-dependent refinements of $\mathcal{S}$ are permitted.
This condition enforces that the internal state retains a unified semantic interpretation across all interventions, rather than serving as a collection of unrelated context-specific representations.
Remarks.
Together, these definitions capture a minimal but stringent notion of representation reuse. They allow for rich internal structure while prohibiting the use of contextual indexing as a representational shortcut. The remainder of this work explores the consequences of these constraints, showing that classical probabilistic models cannot, in general, satisfy them without incurring an irreducible information-theoretic cost.
II.2 Classical Probabilistic Representations
We now specify what we mean by a classical probabilistic representation within the single-state framework introduced above. Throughout, we adopt a representation-level notion of classicality, independent of any specific physical implementation.
Definition 4 (Classical Probabilistic Representation).
A classical probabilistic representation is a single-state representation in which the internal state $s\in\mathcal{S}$ is associated with a classical random variable, and outcome statistics are generated by conditional probability distributions over observable outcomes given $s$. All randomness is assumed to be classical, and joint statistics across interventions are assumed to admit a single underlying probability space.
This definition captures the standard notion of classical simulation: observable statistics arise from probabilistic mixtures over underlying internal states, with no additional contextual degrees of freedom beyond those encoded in $\mathcal{S}$.
Crucially, under the single-state constraint, the same internal state space $\mathcal{S}$ must support all interventions without contextual indexing. Any dependence on context must therefore be mediated through the action of interventions on $\mathcal{S}$, rather than through context-dependent refinement of the underlying probability space.
Definition 5 (Context Information).
We define the context information as the amount of information required to account for contextual dependence in observable statistics beyond what can be encoded in the internal state $s\in\mathcal{S}$. This quantity measures the additional representational resources needed to reconcile multiple interventions within a single classical probability space.
If a classical probabilistic representation could faithfully encode all contextual dependence within the internal state alone, the required context information would vanish. However, as we show below, this is generically impossible under the single-state constraint.
The central question addressed in the remainder of this work can now be stated precisely: Can a classical probabilistic representation, constrained to a fixed internal state space reused across interventions, account for contextual dependence without incurring additional information-theoretic cost? In the next section, we show that the answer is negative. Specifically, we prove that any such representation necessarily requires an irreducible excess of contextual information, establishing a fundamental obstruction to classical single-state representations.
Operational random variables. To make information-theoretic statements explicit, we introduce the following random variables. Let $C$ denote the intervention (context) selected from a set $\mathcal{C}$, $S\in\mathcal{S}$ the internal state, and $O$ the observable outcome. A classical single-state model specifies conditional distributions $p(o\mid s,c)$ together with a prior over $S$. To represent contextual dependence without indexing $\mathcal{S}$ by $c$, the model may require additional contextual information $M$ (an external label or memory register) that is not contained in $S$. All entropies and mutual informations are Shannon quantities [3].
III Information-Theoretic Obstruction
III.1 Main Theorem
We now state the central result of this work. It establishes an unavoidable information-theoretic obstruction faced by classical probabilistic representations that maintain a fixed internal state space across multiple interventions.
Theorem 1 (Information-Theoretic Obstruction to Single-State Classical Representations).
Let $C$ denote the intervention, $S$ the internal state, and $O$ the outcome. Consider any classical single-state representation that reproduces the observed family of conditionals $\{p(o\mid c)\}_{c\in\mathcal{C}}$ via a single underlying probability space. If the observed statistics exhibit contextual dependence across interventions, then any such representation requires an additional contextual variable $M$ (not contained in $S$) such that
$H(M)\;\geq\;I(C;O\mid S)\;>\;0.$ (1)
Equivalently, the required contextual information strictly exceeds what can be attributed to the internal state alone. We write this irreducible excess in the shorthand form $H>I$ in the main text.
Operationally, $H(M)$ quantifies the additional contextual information (in Shannon bits, on average) required to reconcile contextual outcome statistics while preserving a single internal state.
III.2 Interpretation and Scope
The inequality in Eq. (1) is not remarkable by its mathematical form. Its significance lies in what it quantifies: an unavoidable representational cost arising from the requirement that a single internal state space be reused coherently across multiple interventions. Importantly, this obstruction does not reflect insufficient state capacity. Even arbitrarily large state spaces cannot eliminate the excess as long as contextual dependence must be supported within a single classical probability space.
The theorem is representation-theoretic and device-independent. It does not assume any particular dynamics, learning procedure, or physical substrate. Rather, it isolates a structural incompatibility between classical probabilistic representations and the requirement of single-state reuse in the presence of contextual dependence.
The proof proceeds by showing that, under the single-state constraint, contextual dependence forces incompatible conditional structures to be represented within a single underlying probability space. Resolving these incompatibilities requires additional information that cannot be encoded in the internal state without introducing context-dependent refinement. A full proof is provided in Appendix A.
IV A Constructive Illustration
The information-theoretic obstruction established in Sec. III is general and representation-theoretic. To show that this obstruction is not merely formal, we briefly introduce a minimal constructive illustration in which contextual dependence arises under the single-state constraint. The purpose of this illustration is purely explanatory: it serves as an existence witness demonstrating that the obstruction corresponds to realizable representational behavior, rather than to an abstract impossibility detached from concrete models.
$c_{1}$$c_{2}$$c_{3}$$p(o_{1},o_{2})$$p(o_{2},o_{3})$$p(o_{3},o_{1})$$\nexists\;p(o_{1},o_{2},o_{3})$
Figure 1: Schematic illustration of contextuality under the single-state constraint. Each pair of interventions admits a consistent marginal distribution, yet no global joint distribution exists that reproduces all pairwise marginals simultaneously. This incompatibility forces any classical single-state representation to introduce additional contextual information.
We consider a system with a fixed internal state space $\mathcal{S}$ that is reused across multiple interventions. Each intervention acts on the same internal state space and produces observable outcomes through intervention-dependent response functions, in accordance with the single-state and consistency conditions introduced in Sec. II. Crucially, no intervention identity is explicitly encoded within $\mathcal{S}$, and no context-dependent refinement or branching of the state space is permitted. Contextual variation is implemented solely through how interventions probe or transform the shared internal state.
As a minimal illustration, consider three interventions $c\in\{1,2,3\}$ with binary outcomes $o\in\{0,1\}$ such that each pair of interventions admits a consistent marginal distribution, while no single joint distribution over all three outcomes exists. Under the single-state constraint, reproducing these statistics requires additional contextual information specifying which incompatible marginal structure is realized, thereby instantiating the information-theoretic obstruction in a concrete setting.
For instance, let $p(o_{i},o_{j}\mid c_{i},c_{j})$ denote consistent pairwise marginals for each $(i,j)$, while no joint distribution $p(o_{1},o_{2},o_{3})$ exists that reproduces all marginals simultaneously. Any classical realization preserving a single internal state must therefore encode which marginal structure is active, requiring additional contextual information. Such marginal compatibility without global consistency is a standard signature of contextuality.
To make this explicit, consider a deterministic internal state $S\in\{0,1\}$ with $P(S=0)=P(S=1)=1/2$. Let the outcome be given by
$O=S\oplus f(C),$
where $f(C)\in\{0,1\}$ is an intervention-dependent bit. Here we assume that $S$ and $C$ are statistically independent, so that any dependence of $O$ on $C$ is not mediated through the internal state. Then $I(S;O)=0$, while
$I(C;O\mid S)=H(f(C))>0,$
since knowledge of $C$ is required to predict $O$ even when $S$ is known. This explicitly demonstrates that contextual dependence enforces $I(C;O\mid S)>0$, rendering the bound in Theorem 1 nontrivial.
Within this setting, one finds that although each intervention admits a well-defined probabilistic description in isolation, the resulting statistics across interventions cannot, in general, be embedded into a single classical probabilistic representation that preserves single-state reuse. The incompatibility does not arise from insufficient state capacity or from specific dynamical assumptions, but from the requirement that a single internal representation support multiple, mutually incompatible statistical dependencies. As a result, the observable behavior exhibits nonclassical statistical structure directly traceable to representational constraints.
This constructive behavior aligns with the general obstruction proven in Sec. III. Any attempt to classically simulate such behavior while maintaining a fixed internal state space necessarily incurs additional contextual information, consistent with the inequality $H>I$. Importantly, the illustration introduces no assumptions beyond those already formalized in the representational framework: it neither presupposes quantum dynamics nor invokes Hilbert space structure. Its sole role is to clarify how contextual dependence can arise operationally under single-state constraints and to provide intuition for the general information-theoretic result.
V Why Nonclassical Probability Appears
Nonclassical probabilistic frameworks circumvent this limitation by relaxing the requirement of a single global probability space. Instead, they allow probabilistic assignments that remain consistent within each intervention while preserving coherence at the level of the internal state. This structural flexibility enables contextual dependence to be represented without introducing external labels or hidden contextual memory, thereby avoiding the information-theoretic excess quantified by the inequality $H>I$.
From a representational perspective, the obstruction arises from the requirement that all observable statistics be embedded into a single global probability space. Nonclassical probabilistic frameworks relax precisely this requirement: probabilities are assigned consistently within each intervention, while global joint representability is replaced by a weaker consistency condition tied to the shared state description. This structural relaxation is sufficient to avoid encoding intervention identity as an explicit variable, without committing to any particular physical interpretation.
In a quantum representation, the state $\rho$ is single and fixed, but outcome probabilities are given by $p(o\mid c)=\mathrm{Tr}(\rho E_{o}^{(c)})$. Crucially, no joint probability distribution over outcomes for different interventions is assumed to exist. As a result, the classical requirement that all conditionals $p(o\mid c)$ arise from a single underlying random variable $S$ is absent. The quantity $I(C;O\mid S)$ is therefore not operationally meaningful in the same way, and no auxiliary contextual variable $M$ is required. From this perspective, quantum probability avoids the obstruction not by encoding context more efficiently, but by relaxing the assumption of global joint representability that underlies classical single-state models.
This viewpoint explains why nonclassical probability appears repeatedly in systems exhibiting contextual dependence, bounded memory, and order effects. Rather than signaling a departure from classical realism or physical intuition, it reflects a principled adjustment of representational structure necessitated by fixed internal state space across interventions.
VI Implications for Adaptive Intelligence
The results of this work have direct implications for adaptive and intelligent systems operating under internal resource constraints. Many such systems must interact with their environment across multiple contexts while reusing a fixed internal state space, due to limitations on memory, representational capacity, or physical degrees of freedom. Our results show that, under these conditions, contextual dependence cannot be represented classically without incurring an irreducible information-theoretic cost.
From this perspective, phenomena such as context sensitivity, order effects, and apparent violations of classical probabilistic reasoning need not be viewed as anomalies or implementation flaws. Instead, they arise naturally as consequences of representational constraints imposed by single-state reuse. Classical representations fail not because they are insufficiently expressive, but because they require additional contextual information that cannot be accommodated within a fixed internal state space.
More broadly, our analysis suggests that nonclassical probabilistic frameworks provide a principled means of representing contextual dependence in bounded adaptive systems. Their relevance lies not in any specific physical interpretation, but in their ability to maintain internal consistency across interventions without expanding representational resources. In this sense, contextuality reflects a fundamental trade-off between adaptability and representational economy, highlighting a general constraint on the design and analysis of intelligent systems.
VII Discussion and Outlook
We have shown that maintaining a fixed internal state space across multiple interventions entails an irreducible information-theoretic cost for classical probabilistic representations whenever contextual dependence is present. This result reframes contextuality as a representational constraint rather than a peculiarity of specific physical theories. The obstruction we identify is principle-level: it is independent of task details, dynamics, learning rules, or physical implementation.
Several clarifications are in order. First, our results do not imply that classical models are universally inadequate, nor that contextual dependence cannot be handled by classical means in general. Rather, the obstruction arises specifically under the requirement of single-state reuse without context-dependent refinement. Relaxing this requirement—by allowing explicit context encoding, additional memory, or intervention-indexed state spaces—removes the obstruction at the cost of increased representational resources. Our analysis thus makes explicit a trade-off that is often left implicit in classical modeling.
Second, while nonclassical probabilistic frameworks naturally accommodate contextual dependence under single-state constraints, our results do not mandate any particular formalism. Quantum probability appears here as one consistent realization of a broader class of nonclassical representations, distinguished not by physical assumptions but by representational structure. Identifying the minimal mathematical ingredients required for such representations remains an open direction for future work.
Finally, we emphasize that an intervention is not assumed to store its identity within the internal state, nor to leave any persistent record accessible to subsequent interventions. Operationally, interventions are memoryless transformations: once the outcome is produced, no record of the applied intervention is retained in the system. This assumption parallels standard physical idealizations such as projective measurement followed by state update, or controlled operations with subsequent reset. Allowing interventions to leave persistent traces would amount to introducing additional memory resources, thereby explicitly violating the single-state constraint.
Looking ahead, the representational perspective developed here suggests several avenues for further investigation. These include sharpening the relationship between information-theoretic cost and representational complexity, exploring extensions to dynamical and learning settings, and examining how similar obstructions manifest in multi-agent or hierarchical systems. More broadly, our results point toward a unified view of contextuality as a fundamental constraint on representation in bounded adaptive systems, bridging concepts from information theory, probability, and the theory of intelligence.
Related work on information- and memory-cost analyses of simulating contextual correlations has been developed in quantum information theory. Our contribution is not to optimize simulation costs for particular contextuality scenarios, but to isolate a representation-level obstruction under the single-state reuse requirement with interventions-as-operations rather than context labels. In this sense, the present theorem is formulated as a general representational principle, independent of specific inequality families or physical implementations.
Appendix A Proof of Theorem 1
This appendix provides a proof of the information-theoretic obstruction stated in Theorem 1. The proof is representation-theoretic in nature and does not rely on any assumptions about dynamics, learning rules, or physical realization.
Throughout this appendix, we express the abstract shorthand $H>I$ used in the main text in the explicit form
$H(M)\geq I(C;O\mid S),$
where $C$ denotes the intervention, $S$ the internal state, $O$ the observable outcome, and $M$ an auxiliary contextual variable.
A.1 Setup
We consider a classical probabilistic representation with a fixed internal state space $\mathcal{S}$, satisfying the single-state and consistency conditions of Definitions 1–3. Let $S\in\mathcal{S}$ denote the internal state random variable (with realizations $s$), and let $\mathcal{I}=\{I_{1},I_{2},\ldots\}$ denote a set of interventions. Observable outcomes are generated according to $p(o\mid s,I_{k})$.
By assumption, all observable statistics across interventions admit a single underlying classical probability space. Contextual dependence is present when no joint distribution over outcomes exists that is consistent with all observed marginals without reference to the intervention index.
A.2 Contextual Dependence and Incompatible Conditionals
Under the single-state constraint, the same internal state $s$ must support all interventions. Contextual dependence therefore manifests as incompatibility among the families of conditional distributions $\{p(o\mid s,I_{k})\}_{k}$ when viewed as arising from a single underlying random variable $s$.
To reconcile these incompatible conditionals within a classical probability space, additional information specifying how the representation should resolve contextual variation is required. This information cannot be encoded in $s$ itself without introducing intervention-dependent refinement of the state space, which is prohibited by the single-state condition.
A.3 Information-Theoretic Bound
In any classical single-state realization with an auxiliary contextual register $M$, outcomes are generated by a channel $C\to(S,M)\to O$, i.e., $p(o\mid s,m,c)=p(o\mid s,m)$ once $M$ carries all context-dependence not contained in $S$. By the data-processing inequality,
$I(C;O\mid S)\;\leq\;I(C;M\mid S).$ (2)
Moreover, $I(C;M\mid S)\leq H(M)$ holds for any random variables. Hence
$I(C;O\mid S)\;\leq\;H(M).$ (3)
Finally, contextual dependence implies $I(C;O\mid S)>0$ for any single-state classical representation that keeps $S$ context-independent; therefore $H(M)>0$ is necessary, yielding Eq. (1). The bound $H(M)\geq I(C;O\mid S)$ is saturated when $M$ carries precisely the minimal information required to distinguish incompatible conditional structures, for example when $M=f(C)$ in the construction of Sec. IV, and no additional correlations with $S$ or $O$ are introduced.
A.4 Independence from Capacity and Dynamics
Importantly, the bound above does not arise from insufficient capacity of the internal state space. Even if $\mathcal{S}$ is arbitrarily large, the requirement that it remain fixed and reused across interventions prevents the absorption of contextual information into $s$ without violating representation consistency. Likewise, the argument does not depend on any specific dynamics or learning procedures, relying solely on representational constraints.
This completes the proof.
Acknowledgements.
This work was supported by SOBIN Institute LLC under Research Grant SP005. The author used ChatGPT5.2 for English editing and takes full responsibility for the content.

### v2

@@
-Contextuality from Single-State Representations: An Information-Theoretic Principle for Adaptive Intelligence
+Contextuality from Single-State Ontological Models: An Information-Theoretic No-Go Theorem
 Abstract
-Adaptive systems often operate across multiple contexts while reusing a fixed internal state space due to constraints on memory, representation, or physical resources.
-Such single-state reuse is ubiquitous in natural and artificial intelligence, yet its fundamental representational consequences remain poorly understood.
-We show that contextuality is not a peculiarity of quantum mechanics, but an inevitable consequence of single-state reuse in classical probabilistic representations.
-Modeling contexts as interventions acting on a shared internal state, we prove that any classical model reproducing contextual outcome statistics must incur an irreducible information-theoretic cost: dependence on context cannot be mediated solely through the internal state.
-We provide a minimal constructive example that explicitly realizes this cost and clarifies its operational meaning.
-We further explain how nonclassical probabilistic frameworks avoid this obstruction by relaxing the assumption of a single global joint probability space, without invoking quantum dynamics or Hilbert space structure.
-Our results identify contextuality as a general representational constraint on adaptive intelligence, independent of physical implementation.
+Contextuality is a central feature of quantum theory, traditionally understood as the impossibility of reproducing quantum measurement statistics using noncontextual ontological models.
+We consider classical ontological models constrained to reuse a single ontic state space across multiple interventions.
+We prove an information-theoretic no-go theorem showing that such models must incur an irreducible contextual information cost: contextual dependence cannot be fully mediated through the ontic state alone and requires additional contextual information beyond it.
+We provide a constructive example illustrating this obstruction and show that it arises solely from the requirement of ontic state reuse within a classical probability space.
+We further explain how quantum theory avoids this obstruction by relaxing the assumption that all measurement statistics arise from a single underlying classical ontic variable.
+These results identify contextuality as a fundamental information-theoretic constraint on classical ontological models and clarify its origin as a limitation on classical representations.
 I Introduction
-Adaptive systems—biological or artificial—must routinely operate across multiple contexts while relying on limited internal resources.
-In many realistic settings, such systems are constrained to maintain a single internal state space that is reused across different interactions, measurements, or decision scenarios.
-Here, “single-state” does not refer to a scalar or binary variable, but to a fixed internal state space—potentially high-dimensional and carrying many bits of information—that cannot be duplicated, indexed, or branched according to context.
-Such constraints arise naturally in perceptual systems with limited working memory, decision-making agents with bounded internal representations, and physical devices whose internal degrees of freedom must be reused across contexts.
-Related constraints have long been discussed across cognitive science and artificial intelligence under notions such as bounded rationality, resource-limited agents, and fixed internal representational architectures (e.g., Herbert A.
-Simon; Models of Bounded Rationality [4]; John R.
-Anderson [1]).
-Despite their ubiquity, the representational consequences of maintaining a fixed internal state space across contexts remain poorly understood.
-A central difficulty lies in how contextual dependence is represented under such constraints.
-Standard classical approaches typically address contextuality by introducing external labels, hidden variables, or context-dependent internal states, effectively enlarging the representational space.
-While successful in many applications, these strategies implicitly relax the single-state constraint by allowing contextual information to be stored or accessed explicitly.
-In contrast, many adaptive systems—such as perceptual agents responding to sequential stimuli under limited working memory or decision systems operating under strict memory reuse—must accommodate contextual changes through interventions that modify information processing, rather than through explicit encoding of which context is active (e.g., Working Memory [2]).
-This distinction motivates a fundamental question: what representational and informational costs are unavoidable when a system is required to maintain a fixed internal state space across multiple interventions?
-In this work, we address this question at the level of representation rather than dynamics or physical implementation.
-We consider systems whose internal state space may be high-dimensional, but must remain fixed and internally consistent across contexts.
-Within this minimal framework, we show that maintaining such representations necessarily entails an irreducible information-theoretic cost when contextual dependence is present.
-Crucially, this cost does not arise from insufficient state capacity, but from the requirement that a single internal state space be reused coherently across multiple interventions.
-This perspective reframes contextuality not as a peculiarity of particular physical theories, but as a general and unavoidable constraint on representation.
-Our approach makes three complementary contributions.
-First, we introduce a minimal operational formulation of fixed-state representations that cleanly separates contextual effects from assumptions about dynamics or physical realization.
-Second, we demonstrate that any classical probabilistic representation satisfying this formulation necessarily incurs an irreducible information-theoretic cost when contextual dependence is present, reflecting a fundamental obstruction to representing multiple interventions within a single internally consistent state space.
-Third, we present a simple constructive mechanism showing how nonclassical statistical structure can emerge directly from representational constraints, without invoking quantum postulates or Hilbert space assumptions.
-Together, these results establish contextuality as a representational principle, rather than a feature tied to specific physical theories.
+Contextuality is a central feature of quantum theory, reflecting the impossibility of reproducing quantum measurement statistics using noncontextual ontological models [4, 9, 7, 12, 8, 1, 5].
+Within the ontological models framework, measurement outcomes are assumed to arise from an underlying ontic state, together with response functions that map ontic states to observable outcomes.
+Contextuality arises when no single ontic state space can consistently account for outcome statistics across multiple measurement contexts.
+Existing no-go theorems establish contextuality as a fundamental constraint on classical ontological models capable of reproducing quantum predictions.
+However, the origin of this constraint from an information-theoretic perspective remains incompletely understood.
+In particular, it is unclear whether contextuality can be characterized as a fundamental limitation on classical representations arising from constraints on information reuse within a fixed ontic state space.
+In this work, we consider classical ontological models constrained to reuse a single ontic state space across multiple interventions.
+Here, “single-state” does not refer to a scalar or binary variable, but to a fixed ontic state space—potentially high-dimensional and carrying arbitrarily large amounts of information—that cannot be duplicated, indexed, or refined according to context.
+All contextual dependence must therefore be mediated through response functions or interventions acting on a common ontic state space, rather than through context-dependent enlargement or branching of the state space itself.
+A central difficulty arises when contextual dependence must be represented under this constraint.
+Standard classical representations resolve contextuality by introducing context-dependent hidden variables or by refining the ontic state space to encode contextual information explicitly.
+Such constructions effectively relax the requirement of single-state reuse by allowing contextual indexing at the ontological level.
+This raises a fundamental question: can contextual dependence be represented within a classical ontological model constrained to reuse a single ontic state space without introducing additional contextual information beyond the ontic state itself?
+We answer this question in the negative.
+Specifically, we establish an information-theoretic no-go theorem showing that any classical ontological model constrained to reuse a single ontic state space across multiple interventions must incur an irreducible contextual information cost.
+This cost is quantified by the conditional mutual information
+$I(C;O\mid\lambda)>0,$
+where $C$ denotes the intervention, $\lambda$ the ontic state, and $O$ the observable outcome.
+This inequality shows that contextual dependence cannot be fully mediated through the ontic state alone, and requires additional contextual information not contained in the ontic state.
+We further provide a constructive example demonstrating this obstruction explicitly, and show how nonclassical probabilistic frameworks avoid this constraint by relaxing the assumption of a single global classical probability space.
+In particular, quantum theory naturally avoids this obstruction because measurement outcomes are not assumed to arise from a single underlying classical ontic variable reproducing all measurement contexts simultaneously.
+Our results place contextuality within a general information-theoretic framework and clarify its role as a fundamental representational constraint on classical ontological models.
+This perspective complements existing no-go theorems by identifying an explicit information-theoretic obstruction associated with ontic state reuse.
+Our result differs fundamentally from existing no-go theorems such as the Kochen–Specker theorem and related contextuality results [9, 12, 8].
+Those results establish the impossibility of reproducing quantum measurement statistics using noncontextual ontological models satisfying functional or measurement noncontextuality assumptions.
+In contrast, the present theorem does not assume any specific algebraic structure of observables or quantum formalism.
+Instead, it identifies an information-theoretic obstruction arising solely from the requirement that a single ontic state space be reused across multiple interventions within a classical probabilistic representation.
+This establishes contextuality as a representational constraint rooted in information reuse, rather than in the specific structure of quantum observables.
 The remainder of this paper is organized as follows.
 In Sec.
-II, we introduce a representational framework based on single-state reuse and contextual interventions, and define the class of classical probabilistic representations considered throughout.
+II, we introduce the formal representational framework and define classical single-state ontological models.
 In Sec.
-III, we establish the central information-theoretic obstruction, proving that classical single-state representations necessarily incur an irreducible contextual information cost under contextual dependence.
+III, we prove the main information-theoretic no-go theorem.
 Sec.
-IV provides a minimal constructive illustration showing that this obstruction corresponds to realizable representational behavior rather than a purely formal limitation.
+IV provides a constructive illustration of the obstruction.
 In Sec.
-V, we explain why nonclassical probabilistic frameworks naturally accommodate single-state reuse without additional information cost.
+V, we explain how quantum probabilistic structure avoids this constraint.
 Sec.
-VI discusses the implications of these results for adaptive and intelligent systems operating under resource constraints, and Sec.
-VII concludes with a discussion of limitations and future directions.
+VI discusses broader implications and connections.
+Sec.
+VII concludes.
 II Representational Framework
-II.1 Single-State Representations and Interventions
+II.1 Ontological Models
+We begin by recalling the standard ontological models framework [12, 8]. An ontological model provides an underlying classical description of operational statistics in terms of an ontic state.
+Definition 1 (Ontological Model).
+An ontological model consists of the following components:
+  1. 1.
+    An ontic state space $\Lambda$, whose elements $\lambda\in\Lambda$ represent underlying physical states.
+  2. 2.
+    Preparation distributions $\mu(\lambda\mid P)$ over ontic states for each preparation procedure $P$.
+  3. 3.
+    Response functions $\xi(o\mid M,\lambda)$ giving the conditional probability of outcome $o$ given measurement $M$ and ontic state $\lambda$.
+Observable statistics are reproduced as
+$p(o\mid P,M)=\int_{\Lambda}\mu(\lambda\mid P)\,\xi(o\mid M,\lambda)\,d\lambda.$ (1)
+All randomness in the model arises from classical probability distributions over ontic states and response functions.
+II.2 Single-State Ontological Models and Interventions
 We begin by formalizing the representational framework considered throughout this work.
 Our goal is to isolate constraints on representation itself, independently of dynamics, learning rules, or physical implementation.
-Definition 1 (Single-State Representation).
-A single-state representation consists of a fixed internal state space $\mathcal{S}$ together with a family of response functions that map elements of $\mathcal{S}$ to observable outcomes.
-The state space $\mathcal{S}$ may be high-dimensional and encode many bits of information, but it is not indexed, duplicated, or partitioned according to context.
-All contextual variation must therefore be mediated through transformations acting on $\mathcal{S}$, rather than through context-dependent enlargement or branching of the state space.
-Importantly, the single-state condition does not impose a bound on state capacity.
-Instead, it constrains how internal information may be reused across contexts: the same underlying state space must support all contextual interactions.
+Definition 2 (Single-State Ontological Model).
+A single-state ontological model is an ontological model satisfying the following conditions:
+  1. 1.
+    The ontic state space $\Lambda$ is fixed and reused across all interventions or measurement contexts.
+  2. 2.
+    The ontic state space $\Lambda$ is not indexed, duplicated, or refined according to the intervention $C$.
+  3. 3.
+    All observable statistics are generated from a single underlying classical probability space over $\Lambda$.
+Under these conditions, all contextual dependence must be mediated through the response functions $\xi(o\mid C,\lambda)$ acting on a common ontic state space, rather than through context-dependent enlargement or branching of $\Lambda$.
 Definition 3 (Interventions).
-A context is modeled as an intervention, defined as a map that acts on the internal state space $\mathcal{S}$ or on the response functions associated with $\mathcal{S}$.
-Interventions modify how the internal state is processed or probed, but do not introduce additional state variables or external context labels accessible to the representation.
-In particular, the identity of the intervention is not assumed to be explicitly encoded within $\mathcal{S}$.
+A context is modeled as an intervention, defined as a map that acts on the ontic state space $\Lambda$ or on the response functions associated with $\Lambda$.
+Interventions modify how the ontic state is processed or probed, but do not introduce additional state variables or external context labels accessible to the representation.
+In particular, the identity of the intervention is not assumed to be explicitly encoded within $\Lambda$.
 This notion distinguishes interventions from classical context labels: while labels index distinct representational resources, interventions act on a common resource that must be reused across contexts.
 Definition 4 (Representation Consistency).
-A single-state representation is said to be consistent if all interventions act on the same internal state space $\mathcal{S}$ and yield outcome statistics that are jointly representable without contradiction.
-Consistency requires that the internal representation does not depend on which interventions are applied before or after, except through their defined action on $\mathcal{S}$.
-In particular, no intervention-specific hidden states or context-dependent refinements of $\mathcal{S}$ are permitted.
-This condition enforces that the internal state retains a unified semantic interpretation across all interventions, rather than serving as a collection of unrelated context-specific representations.
+A single-state ontological model is said to be consistent if all interventions act on the same ontic state space $\Lambda$ and yield outcome statistics that are jointly representable without contradiction.
+Consistency requires that the ontic representation does not depend on which interventions are applied before or after, except through their defined action on $\Lambda$.
+In particular, no intervention-specific hidden states or context-dependent refinements of $\Lambda$ are permitted.
+This condition enforces that the ontic state retains a unified semantic interpretation across all interventions, rather than serving as a collection of unrelated context-specific representations.
 Remarks.
@@
 The remainder of this work explores the consequences of these constraints, showing that classical probabilistic models cannot, in general, satisfy them without incurring an irreducible information-theoretic cost.
-II.2 Classical Probabilistic Representations
-We now specify what we mean by a classical probabilistic representation within the single-state framework introduced above.
-Throughout, we adopt a representation-level notion of classicality, independent of any specific physical implementation.
-Definition 4 (Classical Probabilistic Representation).
-A classical probabilistic representation is a single-state representation in which the internal state $s\in\mathcal{S}$ is associated with a classical random variable, and outcome statistics are generated by conditional probability distributions over observable outcomes given $s$.
-All randomness is assumed to be classical, and joint statistics across interventions are assumed to admit a single underlying probability space.
-This definition captures the standard notion of classical simulation: observable statistics arise from probabilistic mixtures over underlying internal states, with no additional contextual degrees of freedom beyond those encoded in $\mathcal{S}$.
-Crucially, under the single-state constraint, the same internal state space $\mathcal{S}$ must support all interventions without contextual indexing.
-Any dependence on context must therefore be mediated through the action of interventions on $\mathcal{S}$, rather than through context-dependent refinement of the underlying probability space.
+Crucially, under the single-state constraint, the same ontic state space $\Lambda$ must support all interventions without contextual indexing.
+Any dependence on context must therefore be mediated through the action of interventions on $\Lambda$, rather than through context-dependent refinement of the underlying probability space.
 Definition 5 (Context Information).
-We define the context information as the amount of information required to account for contextual dependence in observable statistics beyond what can be encoded in the internal state $s\in\mathcal{S}$.
+We define the context information as the amount of information required to account for contextual dependence in observable statistics beyond what can be encoded in the ontic state $\lambda\in\Lambda$.
 This quantity measures the additional representational resources needed to reconcile multiple interventions within a single classical probability space.
-If a classical probabilistic representation could faithfully encode all contextual dependence within the internal state alone, the required context information would vanish.
+If a classical probabilistic representation could faithfully encode all contextual dependence within the ontic state alone, the required context information would vanish.
 However, as we show below, this is generically impossible under the single-state constraint.
-The central question addressed in the remainder of this work can now be stated precisely: Can a classical probabilistic representation, constrained to a fixed internal state space reused across interventions, account for contextual dependence without incurring additional information-theoretic cost?
+The central question addressed in the remainder of this work can now be stated precisely: Can a classical probabilistic representation, constrained to a fixed ontic state space reused across interventions, account for contextual dependence without incurring additional information-theoretic cost?
 In the next section, we show that the answer is negative.
@@
 Operational random variables.
-To make information-theoretic statements explicit, we introduce the following random variables.
-Let $C$ denote the intervention (context) selected from a set $\mathcal{C}$, $S\in\mathcal{S}$ the internal state, and $O$ the observable outcome.
-A classical single-state model specifies conditional distributions $p(o\mid s,c)$ together with a prior over $S$.
-To represent contextual dependence without indexing $\mathcal{S}$ by $c$, the model may require additional contextual information $M$ (an external label or memory register) that is not contained in $S$.
+Let $C$ denote the intervention (measurement context), $\lambda\in\Lambda$ the ontic state, and $O$ the observable outcome.
+A classical ontological model specifies:
+  1. 1.
+    A preparation distribution $\mu(\lambda)$ over ontic states.
+  2. 2.
+    Response functions $\xi(o\mid C,\lambda)$ giving outcome probabilities conditioned on the ontic state and intervention.
+Observable statistics are given by
+$p(o\mid C)=\int_{\Lambda}\mu(\lambda)\,\xi(o\mid C,\lambda)\,d\lambda.$ (2)
+To represent contextual dependence within a classical single-state ontological model, an auxiliary contextual variable $M$ may be required that is not contained in $\lambda$.
 All entropies and mutual informations are Shannon quantities [10].
-III Information-Theoretic Obstruction
+III Information-Theoretic No-Go Theorem
 III.1 Main Theorem
 We now state the central result of this work.
-It establishes an unavoidable information-theoretic obstruction faced by classical probabilistic representations that maintain a fixed internal state space across multiple interventions.
-Theorem 1 (Information-Theoretic Obstruction to Single-State Classical Representations).
-Let $C$ denote the intervention, $S$ the internal state, and $O$ the outcome.
-Consider any classical single-state representation that reproduces the observed family of conditionals $\{p(o\mid c)\}_{c\in\mathcal{C}}$ via a single underlying probability space.
-If the observed statistics exhibit contextual dependence across interventions, then any such representation requires an additional contextual variable $M$ (not contained in $S$) such that
-$H(M)\;\geq\;I(C;O\mid S)\;>\;0.$ (1)
-Equivalently, the required contextual information strictly exceeds what can be attributed to the internal state alone. We write this irreducible excess in the shorthand form $H>I$ in the main text.
-Operationally, $H(M)$ quantifies the additional contextual information (in Shannon bits, on average) required to reconcile contextual outcome statistics while preserving a single internal state.
+It establishes an unavoidable information-theoretic obstruction faced by classical probabilistic representations that maintain a fixed ontic state space across multiple interventions.
+Theorem 1 (Information-Theoretic No-Go Theorem for Single-State Ontological Models).
+Let $C$ denote the intervention, $\lambda\in\Lambda$ the ontic state, and $O$ the observable outcome.
+Consider any classical ontological model satisfying:
+  1. 1.
+    A single ontic state space $\Lambda$ is reused across all interventions.
+  2. 2.
+    The ontic state space $\Lambda$ is not indexed or refined according to the intervention $C$.
+  3. 3.
+    Outcome statistics arise from a single underlying classical probability space over $\Lambda$.
+If the observable statistics exhibit contextual dependence, then any such ontological model requires an additional contextual variable $M$, not contained in $\lambda$, such that
+$H(M)\;\geq\;I(C;O\mid\lambda)\;>\;0.$ (3)
+Equivalently, contextual dependence cannot be fully mediated through the ontic state alone.
+This theorem establishes a fundamental information-theoretic obstruction to classical ontological models constrained by ontic state reuse.
+The quantity $I(C;O\mid\lambda)$ quantifies the contextual dependence that cannot be encoded in the ontic state alone.
+The inequality therefore shows that any classical ontological model reproducing contextual statistics must introduce additional contextual information beyond the ontic state, establishing a no-go result for single-state ontological representations.
 III.2 Interpretation and Scope
@@
 (3) is not remarkable by its mathematical form.
-Its significance lies in what it quantifies: an unavoidable representational cost arising from the requirement that a single internal state space be reused coherently across multiple interventions.
+Its significance lies in what it quantifies: an unavoidable representational cost arising from the requirement that a single ontic state space be reused coherently across multiple interventions.
 Importantly, this obstruction does not reflect insufficient state capacity.
@@
 The proof proceeds by showing that, under the single-state constraint, contextual dependence forces incompatible conditional structures to be represented within a single underlying probability space.
-Resolving these incompatibilities requires additional information that cannot be encoded in the internal state without introducing context-dependent refinement.
+Resolving these incompatibilities requires additional information that cannot be encoded in the ontic state without introducing context-dependent refinement.
 A full proof is provided in Appendix A.
 IV A Constructive Illustration
-The information-theoretic obstruction established in Sec.
-III is general and representation-theoretic.
+The information-theoretic no-go theorem established in Sec.
+III identifies a general and representation-theoretic obstruction.
 To show that this obstruction is not merely formal, we briefly introduce a minimal constructive illustration in which contextual dependence arises under the single-state constraint.
@@
 $c_{1}$$c_{2}$$c_{3}$$p(o_{1},o_{2})$$p(o_{2},o_{3})$$p(o_{3},o_{1})$$\nexists\;p(o_{1},o_{2},o_{3})$
-Figure 1: Schematic illustration of contextuality under the single-state constraint.
-Each pair of interventions admits a consistent marginal distribution, yet no global joint distribution exists that reproduces all pairwise marginals simultaneously.
-This incompatibility forces any classical single-state representation to introduce additional contextual information.
-We consider a system with a fixed internal state space $\mathcal{S}$ that is reused across multiple interventions.
-Each intervention acts on the same internal state space and produces observable outcomes through intervention-dependent response functions, in accordance with the single-state and consistency conditions introduced in Sec.
+Figure 1: Illustration of contextual behavior arising under the single-state ontological constraint.
+Each pair of interventions admits a consistent marginal distribution, yet no global joint distribution over outcomes exists that reproduces all pairwise marginals simultaneously.
+This incompatibility demonstrates the information-theoretic obstruction of Theorem 1, showing that any classical ontological model constrained to reuse a fixed ontic state space $\Lambda$ must introduce additional contextual information.
+We consider a system with a fixed ontic state space $\Lambda$ that is reused across multiple interventions.
+Each intervention acts on the same ontic state space and produces observable outcomes through intervention-dependent response functions, in accordance with the single-state and consistency conditions introduced in Sec.
 II.
-Crucially, no intervention identity is explicitly encoded within $\mathcal{S}$, and no context-dependent refinement or branching of the state space is permitted.
-Contextual variation is implemented solely through how interventions probe or transform the shared internal state.
+Crucially, no intervention identity is explicitly encoded within $\Lambda$, and no context-dependent refinement or branching of the state space is permitted.
+Contextual variation is implemented solely through how interventions probe or transform the shared ontic state.
 As a minimal illustration, consider three interventions $c\in\{1,2,3\}$ with binary outcomes $o\in\{0,1\}$ such that each pair of interventions admits a consistent marginal distribution, while no single joint distribution over all three outcomes exists.
@@
 For instance, let $p(o_{i},o_{j}\mid c_{i},c_{j})$ denote consistent pairwise marginals for each $(i,j)$, while no joint distribution $p(o_{1},o_{2},o_{3})$ exists that reproduces all marginals simultaneously.
-Any classical realization preserving a single internal state must therefore encode which marginal structure is active, requiring additional contextual information.
+Any classical realization preserving a single ontic state must therefore encode which marginal structure is active, requiring additional contextual information.
 Such marginal compatibility without global consistency is a standard signature of contextuality.
-To make this explicit, consider a deterministic internal state $S\in\{0,1\}$ with $P(S=0)=P(S=1)=1/2$. Let the outcome be given by
-$O=S\oplus f(C),$
+To make this explicit, consider a deterministic ontic state $\lambda\in\{0,1\}$ with $P(\lambda=0)=P(\lambda=1)=1/2$. Let the outcome be given by
+$O=\lambda\oplus f(C),$
 where $f(C)\in\{0,1\}$ is an intervention-dependent bit.
-Here we assume that $S$ and $C$ are statistically independent, so that any dependence of $O$ on $C$ is not mediated through the internal state.
-Then $I(S;O)=0$, while
-$I(C;O\mid S)=H(f(C))>0,$
-since knowledge of $C$ is required to predict $O$ even when $S$ is known.
-This explicitly demonstrates that contextual dependence enforces $I(C;O\mid S)>0$, rendering the bound in Theorem 1 nontrivial.
+Here we assume that $\lambda$ and $C$ are statistically independent, so that any dependence of $O$ on $C$ is not mediated through the ontic state.
+Then $I(\lambda;O)=0$, while
+$I(C;O\mid\lambda)=H(f(C))>0,$
+since knowledge of $C$ is required to predict $O$ even when $\lambda$ is known.
+This explicitly demonstrates that contextual dependence enforces $I(C;O\mid\lambda)>0$, rendering the bound in Theorem 1 nontrivial.
 Within this setting, one finds that although each intervention admits a well-defined probabilistic description in isolation, the resulting statistics across interventions cannot, in general, be embedded into a single classical probabilistic representation that preserves single-state reuse.
@@
 III.
-Any attempt to classically simulate such behavior while maintaining a fixed internal state space necessarily incurs additional contextual information, consistent with the inequality $H>I$.
+Any attempt to classically simulate such behavior while maintaining a fixed ontic state space necessarily incurs additional contextual information, consistent with the bound in Eq.
+(3).
 Importantly, the illustration introduces no assumptions beyond those already formalized in the representational framework: it neither presupposes quantum dynamics nor invokes Hilbert space structure.
 Its sole role is to clarify how contextual dependence can arise operationally under single-state constraints and to provide intuition for the general information-theoretic result.
-V Why Nonclassical Probability Appears
-Nonclassical probabilistic frameworks circumvent this limitation by relaxing the requirement of a single global probability space.
-Instead, they allow probabilistic assignments that remain consistent within each intervention while preserving coherence at the level of the internal state.
-This structural flexibility enables contextual dependence to be represented without introducing external labels or hidden contextual memory, thereby avoiding the information-theoretic excess quantified by the inequality $H>I$.
-From a representational perspective, the obstruction arises from the requirement that all observable statistics be embedded into a single global probability space.
-Nonclassical probabilistic frameworks relax precisely this requirement: probabilities are assigned consistently within each intervention, while global joint representability is replaced by a weaker consistency condition tied to the shared state description.
-This structural relaxation is sufficient to avoid encoding intervention identity as an explicit variable, without committing to any particular physical interpretation.
-In a quantum representation, the state $\rho$ is single and fixed, but outcome probabilities are given by $p(o\mid c)=\mathrm{Tr}(\rho E_{o}^{(c)})$.
-Crucially, no joint probability distribution over outcomes for different interventions is assumed to exist.
-As a result, the classical requirement that all conditionals $p(o\mid c)$ arise from a single underlying random variable $S$ is absent.
-The quantity $I(C;O\mid S)$ is therefore not operationally meaningful in the same way, and no auxiliary contextual variable $M$ is required.
-From this perspective, quantum probability avoids the obstruction not by encoding context more efficiently, but by relaxing the assumption of global joint representability that underlies classical single-state models.
-This viewpoint explains why nonclassical probability appears repeatedly in systems exhibiting contextual dependence, bounded memory, and order effects.
-Rather than signaling a departure from classical realism or physical intuition, it reflects a principled adjustment of representational structure necessitated by fixed internal state space across interventions.
+V Why Quantum Theory Avoids the Obstruction
+The information-theoretic no-go theorem established in Sec.
+III applies to classical ontological models in which all outcome statistics arise from a single underlying classical probability space over an ontic state space $\Lambda$.
+We now explain why quantum theory avoids this obstruction.
+In an ontological model, observable statistics are generated according to
+$p(o\mid C)=\int_{\Lambda}\mu(\lambda)\,\xi(o\mid C,\lambda)\,d\lambda,$ (4)
+where $\lambda\in\Lambda$ is the ontic state, $\mu(\lambda)$ the preparation distribution, and $\xi(o\mid C,\lambda)$ the response function.
+The no-go theorem shows that if a single ontic state space $\Lambda$ is reused across multiple interventions, and all outcome statistics arise from a single classical probability space, then contextual dependence requires additional contextual information beyond $\lambda$.
+Quantum theory does not satisfy these assumptions. Instead, observable statistics are generated according to the Born rule,
+$p(o\mid C)=\mathrm{Tr}\big(\rho\,E_{o}^{(C)}\big),$ (5)
+where $\rho$ is the quantum state and $\{E_{o}^{(C)}\}$ is the positive operator-valued measure corresponding to intervention $C$.
+Crucially, quantum theory does not assume that all measurement outcomes arise from a single underlying classical ontic variable reproducing all measurement contexts simultaneously.
+Indeed, the Kochen–Specker theorem shows that no noncontextual ontological model exists in which measurement outcomes are determined by a single underlying classical ontic state [9, 12, 8, 1].
+From an information-theoretic perspective, the obstruction identified in Theorem 1 arises from the requirement that contextual dependence be represented within a single classical probability space over ontic states.
+Quantum theory avoids this obstruction by relaxing this requirement.
+While a single quantum state $\rho$ is reused across measurement contexts, measurement outcomes are not assumed to arise from a single underlying classical random variable reproducing all measurement statistics simultaneously.
+Equivalently, quantum probability does not admit a representation in which all measurement outcome statistics are embedded into a single global classical probability space over ontic states without contextual refinement.
+As a result, the conditional mutual information $I(C;O\mid\lambda)$ does not arise in the same manner, and no auxiliary contextual variable is required to mediate contextual dependence.
+From this perspective, quantum theory avoids the information-theoretic obstruction not by encoding contextual information more efficiently, but by relaxing the assumption that observable statistics arise from a single underlying classical ontic state space.
+This structural difference explains why contextuality appears naturally in quantum theory and highlights its origin as a fundamental limitation on classical ontological representations.
 VI Implications for Adaptive Intelligence
+We emphasize that the following discussion is interpretive and does not enter the proof of the main no-go theorem.
 The results of this work have direct implications for adaptive and intelligent systems operating under internal resource constraints.
-Many such systems must interact with their environment across multiple contexts while reusing a fixed internal state space, due to limitations on memory, representational capacity, or physical degrees of freedom.
+Many such systems must interact with their environment across multiple contexts while reusing a fixed ontic state space, due to limitations on memory, representational capacity, or physical degrees of freedom.
 Our results show that, under these conditions, contextual dependence cannot be represented classically without incurring an irreducible information-theoretic cost.
@@
 Instead, they arise naturally as consequences of representational constraints imposed by single-state reuse.
-Classical representations fail not because they are insufficiently expressive, but because they require additional contextual information that cannot be accommodated within a fixed internal state space.
+Classical representations fail not because they are insufficiently expressive, but because they require additional contextual information that cannot be accommodated within a fixed ontic state space.
 More broadly, our analysis suggests that nonclassical probabilistic frameworks provide a principled means of representing contextual dependence in bounded adaptive systems.
@@
 VII Discussion and Outlook
-We have shown that maintaining a fixed internal state space across multiple interventions entails an irreducible information-theoretic cost for classical probabilistic representations whenever contextual dependence is present.
-This result reframes contextuality as a representational constraint rather than a peculiarity of specific physical theories.
-The obstruction we identify is principle-level: it is independent of task details, dynamics, learning rules, or physical implementation.
-Several clarifications are in order.
-First, our results do not imply that classical models are universally inadequate, nor that contextual dependence cannot be handled by classical means in general.
-Rather, the obstruction arises specifically under the requirement of single-state reuse without context-dependent refinement.
-Relaxing this requirement—by allowing explicit context encoding, additional memory, or intervention-indexed state spaces—removes the obstruction at the cost of increased representational resources.
-Our analysis thus makes explicit a trade-off that is often left implicit in classical modeling.
-Second, while nonclassical probabilistic frameworks naturally accommodate contextual dependence under single-state constraints, our results do not mandate any particular formalism.
-Quantum probability appears here as one consistent realization of a broader class of nonclassical representations, distinguished not by physical assumptions but by representational structure.
-Identifying the minimal mathematical ingredients required for such representations remains an open direction for future work.
-Finally, we emphasize that an intervention is not assumed to store its identity within the internal state, nor to leave any persistent record accessible to subsequent interventions.
-Operationally, interventions are memoryless transformations: once the outcome is produced, no record of the applied intervention is retained in the system.
-This assumption parallels standard physical idealizations such as projective measurement followed by state update, or controlled operations with subsequent reset.
-Allowing interventions to leave persistent traces would amount to introducing additional memory resources, thereby explicitly violating the single-state constraint.
-Looking ahead, the representational perspective developed here suggests several avenues for further investigation.
-These include sharpening the relationship between information-theoretic cost and representational complexity, exploring extensions to dynamical and learning settings, and examining how similar obstructions manifest in multi-agent or hierarchical systems.
-More broadly, our results point toward a unified view of contextuality as a fundamental constraint on representation in bounded adaptive systems, bridging concepts from information theory, probability, and the theory of intelligence.
-Related work on information- and memory-cost analyses of simulating contextual correlations has been developed in quantum information theory.
-Our contribution is not to optimize simulation costs for particular contextuality scenarios, but to isolate a representation-level obstruction under the single-state reuse requirement with interventions-as-operations rather than context labels.
-In this sense, the present theorem is formulated as a general representational principle, independent of specific inequality families or physical implementations.
+We have established an information-theoretic no-go theorem for classical ontological models constrained to reuse a single ontic state space across multiple interventions.
+Specifically, we have shown that contextual dependence necessarily incurs an irreducible contextual information cost quantified by the conditional mutual information $I(C;O\mid\lambda)$.
+This result identifies contextuality as a fundamental representational constraint arising from limitations on information reuse within classical ontological models.
+Our result complements existing no-go theorems such as the Kochen–Specker theorem [9, 12] by identifying an explicit information-theoretic obstruction.
+Rather than arising from algebraic properties of observables or Hilbert space structure, the obstruction emerges directly from the requirement that a single ontic state space be reused across multiple interventions without contextual refinement.
+This provides a new perspective on contextuality as an information-theoretic limitation on classical representations.
+Nonclassical probabilistic frameworks avoid this obstruction by relaxing the assumption of a single global classical probability space.
+In quantum theory, a single quantum state $\rho$ is reused across multiple measurement contexts, but measurement outcomes are not assumed to arise from a single underlying classical ontic variable reproducing all measurement statistics simultaneously.
+This structural feature allows contextual dependence to be represented without introducing auxiliary contextual information external to the quantum state.
+Beyond quantum foundations, similar representational constraints arise naturally in adaptive and resource-limited systems.
+For example, perceptual systems with limited working memory, decision-making agents with bounded internal representations, and physical devices with fixed internal degrees of freedom must often reuse a single ontic state space across multiple interactions.
+Related constraints have been studied extensively in cognitive science and artificial intelligence under notions such as bounded rationality, resource-limited representation, and working memory limitations [11, 2, 3, 5].
+From this perspective, contextuality can be understood more generally as a consequence of representational constraints imposed by ontic state reuse.
+Our analysis is representation-theoretic and does not depend on specific physical dynamics or implementation details.
+The obstruction arises solely from the structural requirement that a single ontic state space be reused across multiple interventions.
+Relaxing this requirement—for example, by allowing context-dependent ontic state refinement or additional contextual memory—removes the obstruction at the cost of increased representational resources.
+Several directions for future work remain.
+These include extending the analysis to dynamical settings, characterizing minimal contextual information costs for specific contextuality scenarios, and exploring connections between contextuality, information complexity, and representational efficiency.
+More broadly, the information-theoretic perspective developed here may provide a useful framework for understanding fundamental limits on classical representations in both physical and abstract systems.
+In summary, our results establish contextuality as a fundamental information-theoretic obstruction to classical ontological models constrained by ontic state reuse.
+This provides a new foundation for understanding contextuality as a general representational constraint arising from limitations on classical probabilistic structure.
 Appendix A Proof of Theorem 1
-This appendix provides a proof of the information-theoretic obstruction stated in Theorem 1.
-The proof is representation-theoretic in nature and does not rely on any assumptions about dynamics, learning rules, or physical realization.
-Throughout this appendix, we express the abstract shorthand $H>I$ used in the main text in the explicit form
-$H(M)\geq I(C;O\mid S),$
-where $C$ denotes the intervention, $S$ the internal state, $O$ the observable outcome, and $M$ an auxiliary contextual variable.
+This appendix provides a proof of Theorem 1.
+The proof is representation-theoretic in nature and does not rely on any assumptions about physical dynamics, learning rules, or implementation details.
+Throughout this appendix, we express the information-theoretic obstruction in explicit form as
+$H(M)\;\geq\;I(C;O\mid\lambda),$ (6)
+where $C$ denotes the intervention (measurement context), $\lambda\in\Lambda$ the ontic state, $O$ the observable outcome, and $M$ an auxiliary contextual variable not contained in the ontic state.
+All entropies and mutual informations are Shannon quantities [10, 6].
 A.1 Setup
-We consider a classical probabilistic representation with a fixed internal state space $\mathcal{S}$, satisfying the single-state and consistency conditions of Definitions 1–3.
-Let $S\in\mathcal{S}$ denote the internal state random variable (with realizations $s$), and let $\mathcal{I}=\{I_{1},I_{2},\ldots\}$ denote a set of interventions.
-Observable outcomes are generated according to $p(o\mid s,I_{k})$.
-By assumption, all observable statistics across interventions admit a single underlying classical probability space.
-Contextual dependence is present when no joint distribution over outcomes exists that is consistent with all observed marginals without reference to the intervention index.
-A.2 Contextual Dependence and Incompatible Conditionals
-Under the single-state constraint, the same internal state $s$ must support all interventions.
-Contextual dependence therefore manifests as incompatibility among the families of conditional distributions $\{p(o\mid s,I_{k})\}_{k}$ when viewed as arising from a single underlying random variable $s$.
-To reconcile these incompatible conditionals within a classical probability space, additional information specifying how the representation should resolve contextual variation is required.
-This information cannot be encoded in $s$ itself without introducing intervention-dependent refinement of the state space, which is prohibited by the single-state condition.
+We consider a classical ontological model with a fixed ontic state space $\Lambda$, satisfying the single-state conditions of Definition 2.
+The ontic state is represented by a random variable $\lambda\in\Lambda$, distributed according to a preparation distribution $\mu(\lambda)$.
+Let $\mathcal{C}$ denote the set of interventions (measurement contexts), with $C\in\mathcal{C}$ a random variable specifying the selected intervention.
+Observable outcomes are generated according to response functions $\xi(o\mid C,\lambda)$, yielding operational statistics
+$p(o\mid C)=\int_{\Lambda}\mu(\lambda)\,\xi(o\mid C,\lambda)\,d\lambda.$ (7)
+By assumption, the ontic state space $\Lambda$ is fixed and reused across all interventions, and all observable statistics arise from a single underlying classical probability space over $\Lambda$.
+Contextual dependence is present when the observable statistics cannot be represented as arising from a context-independent ontic description alone. Formally, this corresponds to the condition
+$I(C;O\mid\lambda)>0.$ (8)
+This inequality expresses that knowledge of the intervention $C$ provides additional information about the outcome $O$ beyond what is contained in the ontic state $\lambda$.
+A.2 Necessity of an Auxiliary Contextual Variable
+Under the single-state constraint, the ontic state space $\Lambda$ is fixed and cannot be refined or indexed by the intervention.
+Therefore, any contextual dependence that cannot be encoded in $\lambda$ must be mediated through an additional contextual variable $M$.
+Operationally, the most general classical ontological model reproducing the statistics can be represented by the probabilistic channel
+$C\;\to\;(\lambda,M)\;\to\;O,$ (9)
+such that the outcome distribution satisfies
+$p(o\mid\lambda,M,C)=p(o\mid\lambda,M).$ (10)
+This condition expresses that once the ontic state $\lambda$ and auxiliary contextual variable $M$ are specified, the outcome is independent of the intervention.
 A.3 Information-Theoretic Bound
-In any classical single-state realization with an auxiliary contextual register $M$, outcomes are generated by a channel $C\to(S,M)\to O$, i.e., $p(o\mid s,m,c)=p(o\mid s,m)$ once $M$ carries all context-dependence not contained in $S$.
-By the data-processing inequality,
-$I(C;O\mid S)\;\leq\;I(C;M\mid S).$ (2)
-Moreover, $I(C;M\mid S)\leq H(M)$ holds for any random variables. Hence
-$I(C;O\mid S)\;\leq\;H(M).$ (3)
-Finally, contextual dependence implies $I(C;O\mid S)>0$ for any single-state classical representation that keeps $S$ context-independent; therefore $H(M)>0$ is necessary, yielding Eq.
-(1).
-The bound $H(M)\geq I(C;O\mid S)$ is saturated when $M$ carries precisely the minimal information required to distinguish incompatible conditional structures, for example when $M=f(C)$ in the construction of Sec.
-IV, and no additional correlations with $S$ or $O$ are introduced.
-A.4 Independence from Capacity and Dynamics
-Importantly, the bound above does not arise from insufficient capacity of the internal state space.
-Even if $\mathcal{S}$ is arbitrarily large, the requirement that it remain fixed and reused across interventions prevents the absorption of contextual information into $s$ without violating representation consistency.
-Likewise, the argument does not depend on any specific dynamics or learning procedures, relying solely on representational constraints.
-This completes the proof.
+We now derive the lower bound on the contextual information required.
+From the channel structure
+$C\;\to\;(\lambda,M)\;\to\;O,$ (11)
+the data-processing inequality implies
+$I(C;O\mid\lambda)\;\leq\;I(C;M\mid\lambda).$ (12)
+This inequality expresses that any contextual dependence observed in $O$ must be mediated through the auxiliary contextual variable $M$ once the ontic state $\lambda$ is fixed.
+Next, using the general information-theoretic bound
+$I(C;M\mid\lambda)\;\leq\;H(M),$ (13)
+we obtain
+$I(C;O\mid\lambda)\;\leq\;H(M).$ (14)
+Rearranging yields the fundamental bound
+$H(M)\;\geq\;I(C;O\mid\lambda).$ (15)
+Finally, contextual dependence implies
+$I(C;O\mid\lambda)>0,$ (16)
+since otherwise the outcome would be conditionally independent of the intervention given the ontic state, contradicting contextual dependence.
+Therefore,
+$H(M)>0,$ (17)
+showing that a strictly positive amount of contextual information is required.
+A.4 Minimality and Saturation
+The bound is tight. Equality
+$H(M)=I(C;O\mid\lambda)$ (18)
+is achieved when the auxiliary contextual variable $M$ carries precisely the minimal information required to resolve contextual dependence, without introducing additional correlations.
+An explicit example achieving saturation is given in Sec.
+IV, where $M$ is constructed as a deterministic function of the intervention $C$ encoding the minimal contextual information necessary to reproduce incompatible conditional structures.
+A.5 Independence from Ontic State Capacity and Dynamics
+Importantly, the bound does not arise from limitations on the size or capacity of the ontic state space $\Lambda$.
+Even if $\Lambda$ is arbitrarily large, the requirement that it remain fixed across interventions prevents contextual information from being absorbed into the ontic state without violating the single-state condition.
+Likewise, the argument does not depend on any assumptions about physical dynamics, measurement disturbance, or learning procedures.
+The obstruction arises purely from representational constraints imposed by ontic state reuse.
+A.6 Conclusion of the Proof
+We have shown that any classical ontological model constrained to reuse a single ontic state space must introduce an auxiliary contextual variable $M$ satisfying
+$H(M)\;\geq\;I(C;O\mid\lambda)\;>\;0.$ (19)
+This establishes the information-theoretic no-go theorem stated in Theorem 1.
+$\square$
 Acknowledgements.